20 men can cut 30 trees in 4 hours. If 4 men leave the job, how
many trees will be cut in 6 hours?

Answer by ptaylor(2198)


20 men work at the rate of 30/4 trees per hour
So one man works at the rate of 30/(4*20)=3/8 tree per hour
And 20-4=16 men works at the rate of (3/8)*16=6 trees per hour
And in 6 hours, 16 men can cut 6*6=36 trees
Another way to look at this problem:
Let x=number of trees that can be cut by 16 men working 6 hours
It takes 20*4=80 man-hours to cut 30 trees
16 men working 6 hours provides 96 man hours
So we can set up a ratio:
80 man-hours is to 30 trees as 96 man-hours is to x trees, or:
80/30=96/x cross multiply
80x=30*96=2880
x=36

Hope this helps---ptaylor